OpenStudy (anonymous):

\[6^{2}+9x-12\div3x\]

OpenStudy (diyadiya):

Divide each terms by \(3x\) first of all :- \(6^2 = 36 \) 36 + 9x - 12 = 24 + 9x is this clear?

OpenStudy (diyadiya):

Now divide each terms by 3x

OpenStudy (diyadiya):

\[\frac{24}{3x}+ \frac{9x}{36}\]Now simplify this.

OpenStudy (anonymous):

ok thank you

OpenStudy (diyadiya):

Welcome :)

OpenStudy (diyadiya):

Did you get the answer?

OpenStudy (anonymous):

no i'm working on it

OpenStudy (diyadiya):

Oops! i made a small typo: it 3x not 36 in the denominator ,sorry! \[\frac{24}{3x}+ \frac{9x}{3x}\]

OpenStudy (anonymous):

l0l its ok